ivan is 5 years older than cyrus. if the sum of their ages is 43 how old is ivan? show solution
Answer by Amily_2190(27)

Cyrus= x
Ivan= x+5 (Ivan is five years older than Cyrus)
x+(x+5)=43 (The sum of their ages is 43)
2x+5=43
2x=38
x=19
Ivan=x+5=19+5=24
Ivan is 24 years old.